OpenStudy (anonymous):

(4x + 3y) - (2x -5y) . simplify. show your work

OpenStudy (anonymous):

4x -2x + 3y + 5y = 2 x + 8y =2 (x + 4y)

OpenStudy (anonymous):

4x + 3y - 2x + 5y =2x + 8y

OpenStudy (aravindg):

4x -2x + 3y + 5y = 2 x + 8y =2 (x + 4y)